OpenStudy (anonymous):

A light house uses a parabolic reflector that is 1 m in diameter how deep should the reflectpr be if the light source is placed halfwayy between the vertex and the plane of the rim?

OpenStudy (mathmale):

If the (general) shape of the reflector is that of a parabola, and if we assume that the parabolic reflector opens UP, then the general form of the equation of that parabola is \[4py=x ^{2}\]

OpenStudy (mathmale):

where p represents the distance between the vertex (0,0) and the focus (0, (d/2)) (Note: d is the unknown depth of the parabolic reflector.) We need to find d, the depth of the parabolic reflector. Note that if the diameter of the reflector at the rim is 1 meter (which is a given), then the radius of the the rim is 1/2 meter. This is an x value. The reflector will be d meters deep, as measured downward from the rim to the x-axis (that is, parallel to the y axis). Thus, starting from \[4py=x ^{2}\] with y=d=depth of reflextor, and x=1/2 meter=radius of rim of reflector, and p=d/2, \[4(\frac{ d }{ 2 })(d)=(\frac{ 1 }{ 2 })^{2}.\] Just solve this for d=the depth of the reflector.

OpenStudy (mathmale):

It may help you to visualize what's happening here if you were to sketch the parabolic reflector on cartesian coordinates and label everything: d=depth of reflector d/2=distance of focus from vertex (0,0) x=1/2 meter = radius of reflector rim reflector diameter at rim = 1 meter